# Mr. DETR: Instructive Multi-Route Training for Detection Transformers

### v3

@@ 4.3 Extension to Instance Segmentation
We integrate our method with the detection transformer architecture for the instance segmentation task. Specifically, following [70, 65], we develop an object mask prediction head based on Deformable-DETR++ [70] utilizing 300 queries. For simplicity, both one-to-one and one-to-many assignments consider only box and classification costs, consistent with the detection transformer. We present the AP metric for both box and mask predictions in Tab. 4. Over a training schedule of 12 and 24 epochs, our auxiliary training method enhances the mask AP by approximately 3.6% and 2.5% compared to the baseline model, respectively. These experimental results highlight the effectiveness of our method.
@@
4.4 Effectiveness Analysis
@@
4.5 Ablation Study
Effectiveness of Multi-route Training.
To evaluate the contribution of each auxiliary route, we perform an ablation study, as shown in Tab. 5. Route-1, with its independent FFN, enhances the baseline model by 2.0% mAP. Experimental results indicate that Route-3 provides a 2.8% improvement on Deformable-DETR++ [70], highlighting the effectiveness of our proposed instructive self-attention. By incorporating multi-route training, Deformable-DETR++ [70] achieves 50.7% mAP, surpassing the baseline by 3.1%.
Route-1 Route-2 Route-3 mAP AP50 AP75
@@ 4.5 Ablation Study
50.7 (+3.1) 68.2 55.4
Table 5: The ablation study of different routes in our method. ‘Route-1’: the auxiliary training route with independent FFN. ‘Route-2’: the primary route for one-to-one prediction. ‘Route-3’: the auxiliary training route with instructive self-attention.
@@
Designs of the Instruction Mechanism. To further reduce the trainable parameters in Route-3 and enhance parameter sharing with the primary route, we introduce the instruction mechanism. In Tab. 6, we evaluate the performance of different designs of the instruction mechanism, as discussed in Sec. 3.3. Initially, we evaluate two conventional variants: one with independent self-attention and another without self-attention in Route-3. Experimental results indicate that our approach surpasses these variants by 0.7% and 0.8% mAP, respectively. Additionally, we implement a variant using independent object queries in Route-3 while sharing all other parameters with the primary route. Our method exceeds this design by 0.8% mAP. We also investigate an implementation that integrates instruction tokens into the object queries by addition, unlike our method that appends them. Experimental results confirm the superiority of our approach, because it dynamically and flexibly guides the object queries to achieve one-to-many predictions.
Variants mAP AP50 AP75
Independent self-attention 50.0 67.7 54.8
Remove self-attention 49.9 67.3 54.7
Independent object queries 49.9 67.5 54.6
Instruction by addition 50.1 67.6 54.7
Instruction by concatenation (ours) 50.7 68.2 55.4
Table 6: Different designs of instructive self-attention.
number 1 5 10 50 100
AP 50.2 50.1 50.7 50.4 50.5
(a) The number of instruction tokens.
layers 0 0$\sim$1 0$\sim$2 0$\sim$3 0$\sim$4 0$\sim$5
AP 50.3 50.2 50.4 50.3 50.4 50.7
(b) Layers of instruction tokens.
Table 7: Ablation study on configurations of instruction tokens.
Impact of Configurations in Instructive Self-Attention. We explore different configurations of instruction tokens. First, we examine the impact of varying the number of instruction tokens, as shown in Tab. 7(a). The results demonstrate that performance improves with an increased number of instruction tokens. Next, we investigate the effect of incorporating instruction tokens across different decoder layers in Tab. 7(b). The experimental results indicate that optimal performance is achieved when all layers utilize instruction tokens. As illustrated in Fig. 6, instruction tokens of all layers significantly contribute to guiding object queries towards one-to-many targets. In our work, we empirically apply instruction tokens across all six decoder layers and set their number to 10.
5 Conclusion
@@
Regarding the model with auxiliary one-to-many training as a multi-task framework, we examine the roles of each component in the transformer decoder for two training targets. Our empirical findings indicate that any independent component in the decoder can effectively learn both one-to-one and one-to-many training targets, even when other components are shared. Building on this insight, we propose a multi-route training mechanism, featuring a primary route and two auxiliary training routes. This mechanism is further enhanced by our novel instructive self-attention, which dynamically and flexibly guides object queries for one-to-many prediction. Importantly, the auxiliary training routes are discarded during inference, ensuring no impact on model architecture or inference cost. Extensive experiments confirm the effectiveness of our method.
@@ Acknowledgement
Appendix
Appendix A Analysis of Instructive Self-Attention
Figure 7: The cosine similarity between 10 instruction tokens used in Deformable-DETR++ [70].
We present the PyTorch-style pseudo-code for our proposed instructive self-attention method in Alg. 1. As detailed in our primary manuscript, our multi-route training approach involves three training routes. Route-2 serves as the primary route designated for one-to-one prediction, which is the same as the baseline model. Route-1 acts as an auxiliary route with an independent FFN aimed at one-to-many predictions. Meanwhile, Route-3 operates as an auxiliary route incorporating our novel instructive self-attention to facilitate one-to-many predictions. In this approach, we establish a collection of trainable tokens as instructions to direct the object queries and subsequent modules in executing one-to-many prediction. These trainable instruction tokens are affixed to the input object queries through concatenation, allowing for the dynamic and adaptable transmission of instructions via self-attention mechanisms. The resulting output from the instruction tokens after undergoing self-attention is not retained, as these tokens do not possess the capability to locate objects.
(a) Evaluation results of box
(b) Evaluation results of mask
Algorithm 1 Pseudo-code of Instructive Self-Attention in a PyTorch-like style.
class InstructAttn(nn.Module):
def __init__(self, embed_dim, num_heads, num_ins):
# embed_dim: the embedding dimension used
# num_heads: the number of heads in self-attention
# num_ins: the number of instruction tokens
self.ins_sa = nn.MultiheadAttention(
embed_dim = embed_dim,
num_heads = num_heads,
# define learnable instruction tokens
self.instruct_o2m = nn.Embedding(num_ins, embed_dim)
self.pos_o2m = nn.Embedding(num_ins, embed_dim)
self.num_ins = num_ins
def forward(self, query, key, value, query_pos, key_pos, route="route-3"):
# route: identify the current route
query = query + query_pos
key = key + key_pos
if route == "route-3":
ins_token = self.instruct_o2m.weight
ins_pos = self.pos_o2m.weight
# concatenate instruction tokens into the input sequence
query = torch.cat([query, ins_token + ins_pos], dim=1)
key = torch.cat([key, ins_token + ins_pos], dim=1)
value = torch.cat([value, ins_token], dim=1)
# perform self-attention
out = self.ins_sa(
query, key, value
if route == "route-3":
# discard the corresponding output of instruction tokens
out = out[:, :-self.num_ins, :]
return out
(a) Ablation of $K$
(b) Ablation of $\alpha$
(c) Ablation of $\tau$
Figure 9: Influence of hyper-parameters $K$, $\alpha$ and $\tau$ in the one-to-many assignment. (a) Influence of $K$ for selecting top-$K$ positive candidates. (b) Influence of $\alpha$ that denotes the weight of classification confidence when forming the matching score $M$. (c) Influence of $\tau$ that is used to filter out low-quality candidates.
In the main paper, we ablate the configurations of the proposed instructive self-attention, including the number of instruction tokens and layers using instruction tokens. Ablation studies demonstrate that the model is not sensitive to the layers and numbers of instruction tokens, due to the following reasons: (i) Regarding the impact of layers of instruction tokens, we hypothesize that the information from instruction tokens in the first layer can be retained and utilized by subsequent layers. Meanwhile, residual connections across transformer decoder layers may help preserve the instruction information for later layers. (ii) To study the impact of the number of instruction tokens, as shown in Fig. 7, we calculate the cosine similarity between the 10 instruction tokens and find that most of them are very similar. This indicates that most instruction tokens may play a similar role, making the model insensitive to their number.
Route Epoch NMS Box Mask
mAP AP50 AP75 APs APm APl mAP AP50 AP75 APs APm APl
Baseline 12 46.5 64.2 50.8 28.8 50.0 60.7 32.4 55.7 32.6 11.9 35.6 54.4
Baseline 12 46.7 65.0 50.7 29.0 50.1 60.9 32.5 56.3 32.6 12.0 35.8 54.8
Route-2 (primary) 12 49.5 66.6 54.1 30.3 52.6 64.7 36.0 59.8 37.2 13.6 39.6 59.7
Route-2 (primary) 12 49.7 67.6 54.1 30.4 52.8 65.0 36.2 60.5 37.2 13.7 39.8 60.1
Route-1 (auxiliary) 12 14.5 18.8 16.0 15.6 21.6 19.1 10.9 17.1 11.6 5.9 13.4 20.2
Route-1 (auxiliary) 12 49.8 67.3 54.6 31.0 53.0 65.3 36.1 60.1 37.2 13.9 39.4 59.9
Route-3 (auxiliary) 12 14.5 18.7 15.9 15.3 22.1 19.7 10.8 17.0 11.5 5.6 13.6 20.3
Route-3 (auxiliary) 12 49.9 67.3 54.6 30.8 53.2 65.4 36.2 60.2 37.3 13.6 39.5 59.8
Baseline 24 48.6 66.4 53.1 30.8 51.9 62.9 35.1 59.1 36.0 14.3 38.9 57.5
Baseline 24 48.6 67.1 52.8 30.9 51.8 63.0 35.2 59.6 35.9 14.4 39.0 57.7
Route-2 (primary) 24 50.3 68.0 54.7 31.5 53.2 65.0 37.6 61.4 38.9 15.1 41.4 60.5
Route-2 (primary) 24 50.4 68.7 54.6 31.6 53.3 65.2 37.7 61.9 38.9 15.1 41.6 60.9
Route-1 (auxiliary) 24 14.0 18.1 15.2 15.4 21.1 18.4 10.7 16.6 11.3 5.8 13.4 19.1
Route-1 (auxiliary) 24 50.4 68.2 55.1 31.5 53.5 65.2 37.6 61.7 38.8 14.9 41.2 60.4
Route-3 (auxiliary) 24 14.1 18.3 15.3 15.6 21.2 18.8 10.7 16.7 11.3 5.9 13.3 19.3
Route-3 (auxiliary) 24 50.4 68.2 54.8 31.7 53.4 65.2 37.6 61.6 38.8 14.9 41.2 60.5
Table 8: The detailed performance of each route in Mr. DETR. The Deformable-DETR++ [70] employing 300 queries serves as our baseline model. ‘Route-2’: the primary route employed for one-to-one prediction, identical in functionality to the baseline model. ‘Route-1’: the auxiliary route for one-to-many prediction, which is built with an independent FFN. ‘Route-3’: the auxiliary route for one-to-many prediction, which utilizes our proposed instructive self-attention.
Appendix B Convergence Curves
Employing Deformable-DETR++ [70] with 300 queries, we perform training on the instance segmentation task both with and without the integration of our proposed approach. Consistent with established methods [70, 61, 65, 20, 21], models are trained using 12 and 24 epoch schedules, respectively. The learning rate is reduced by a factor of 0.1 at the 11th and 20th epochs according to the 12 and 24 epoch schedules, respectively. We illustrate the evaluation results for bounding box predictions in Fig. 8(a) and for instance mask predictions in Fig. 8(b). The evaluation results demonstrate that our approach significantly enhances the training process of the baseline model.
Appendix C Impact of Hyper-Parameters
Fig. 9 shows the influence of the hyper-parameters $K$, $\alpha$, and $\tau$ in the one-to-many assignment [42, 65, 20]. In Fig. 9(a), we observe that as the number of positive candidates increases, the model achieves its highest performance when $K=6$. However, when $K>7$, the one-to-many assignment increases the difficulty of removing duplicates in the primary route, leading to decreased performance. For the weight of classification confidence $\alpha$, as shown in Fig. 9(b), the model achieves the best performance at $\alpha=0.3$. Similarly, in Fig. 9(c), the performance improves as the filter threshold $\tau$ increases, reaching its peak at $\tau=0.4$. Beyond this value, the performance declines, potentially due to the filtering out of many high-quality candidates. In our method, we empirically set $K=6$, $\alpha=0.3$, and $\tau=0.4$.
Appendix D Detailed Performance of Each Route
As detailed in Sec. B, we train the Mr. DETR for the instance segmentation task based on Deformable-DETR++ [70]. We report the evaluation results of each route of our method in Tab. 8. Experimental results indicate that the primary route is adept at accomplishing one-to-one prediction, as evidenced by the fact that Non-Maximum-Suppression (NMS) yields only a minor improvement of approximately 0.1 - 0.2% in both box mAP and mask mAP. Conversely, for the auxiliary routes, the application of NMS significantly improves the performance (about 35% and 25% in terms of box mAP and mask mAP) of Route-1 and Route-3, highlighting their capability for effective one-to-many prediction. This further substantiates that our introduced instructive self-attention is proficient in efficiently guiding object queries for one-to-many prediction.
Models Epochs Queries mAP AP50 AP75 APs APm APl
Deformable-DETR++ [70] 4 900 30.4 40.8 33.1 16.1 30.1 39.1
w/ Mr. DETR 4 900 32.7 (+2.3) 42.7 (+1.9) 35.8 (+2.7) 17.1 (+1.0) 32.3 (+2.2) 42.6 (+3.5)
Table 9: Experiments on the large-scale Obejcts365 dataset [46]. Deformable-DETR++ and our model use the ResNet-50 [17] backbone.
Appendix E Experiments on the Objects365
Objects365 [46] is a large-scale dataset with 365 classes, which contains about 2,000,000 images. To further verify the scalability of our method on the large-scale dataset, we conduct experiments on Objects365 using Deformable-DETR++ [70] model with 900 queries. To save training time, we train the baseline model and our method for 4 epochs only. The initial learning rate is set to 2e-4 and decays at the third epoch. Other training settings are the same as the model trained on the COCO dataset [32]. We report the performance comparison in Tab. 9.
Appendix F Training Cost
We measure the training time of different methods in Tab. 10. The training time denotes the average duration of each epoch. We evaluate training time on 8 NVIDIA 3090 GPUs with a batch size of 16. The experimental results indicate that our proposed method achieves an effective trade-off between performance and training costs.
Model Time GFLOPs mAP
Deformable-DETR++ [70] 84 234.9 47.0
H-DETR [21] 104 (+20) 265.0 48.7 (+1.7)
DAC-DETR [20] 94 (+10) - -
MS-DETR [65] 96 (+12) - 48.8 (+1.8)
Group-DETR [5] 123 (+39) - -
Mr. DETR (Ours) 101 (+17) 258.0 49.5 (+2.5)
Table 10: Comparison of training time (minutes) and training GFLOPs of various methods. All methods utilize Deformable-DETR++ with 300 queries as the baseline. The training time represents the average duration per training epoch.
Appendix G Performance of Intermediate Layers
Typically, DETR-like object detectors consist of six layers each in their transformer encoders and decoders. As mentioned in Sec. B and Sec. D, our approach for mask prediction is based on Deformable-DETR++ [70], utilizing solely the last decoder layer. All six decoder layers are employed for object detection tasks. Therefore, we only evaluate the box prediction for all layers as shown in Tab. 11. Evaluation results suggest that our method can effectively improve the performance of the primary route across all six decoder layers, demonstrating the efficacy of our approach. Moreover, the one-to-many prediction training routes, namely, Route-1 and Route-3, significantly surpass the primary route in the shallower layers. For example, with a model trained on a 12-epoch schedule, Route-3 achieves a 6.4% improvement over the primary route in layer 0, and a 0.4% improvement in layer 5. These experiments indicate that the primary route needs more decoder layers to reach comparable performance as the auxiliary routes equipped with NMS. For instance, Route-1 and Route-3 can reach 49.4% mAP in layer 2, while the primary route achieves 49.4% mAP in layer 4.
Method Layer Route Epoch NMS mAP AP50 AP75 APs APm APl
Baseline [70] 0 - 12 38.5 53.4 42.2 23.1 41.8 50.0
w/ Mr.DETR 0 Route-2 41.0 (+2.5) 55.5 45.0 24.8 43.8 54.3
w/ Mr.DETR 0 Route-1 47.5 64.9 52.2 29.0 51.0 62.1
w/ Mr.DETR 0 Route-3 47.4 64.9 52.3 29.1 50.9 62.4
Baseline [70] 1 - 12 42.6 58.7 46.5 25.7 46.2 55.1
w/ Mr.DETR 1 Route-2 45.8 (+3.2) 61.8 50.3 27.6 49.0 60.3
w/ Mr.DETR 1 Route-1 48.9 66.1 53.5 30.4 52.4 63.8
w/ Mr.DETR 1 Route-3 49.0 66.3 53.6 30.4 52.6 64.2
Baseline [70] 2 - 12 45.0 62.0 49.2 27.7 48.7 58.6
w/ Mr.DETR 2 Route-2 48.0 (+3.0) 64.7 52.5 29.3 51.3 62.4
w/ Mr.DETR 2 Route-1 49.4 66.7 54.2 31.1 52.8 64.3
w/ Mr.DETR 2 Route-3 49.4 66.5 54.1 31.0 52.8 64.3
Baseline [70] 3 - 12 46.1 63.5 50.4 28.6 49.6 60.1
w/ Mr.DETR 3 Route-2 49.0 (+2.9) 66.0 53.5 30.4 52.3 63.7
w/ Mr.DETR 3 Route-1 49.7 67.1 54.5 31.3 53.0 64.8
w/ Mr.DETR 3 Route-3 49.9 67.2 54.6 31.4 53.3 64.8
Baseline [70] 4 - 12 46.4 64.1 50.7 29.0 50.0 60.7
w/ Mr.DETR 4 Route-2 49.4 (+3.0) 66.5 54.0 30.4 52.6 64.3
w/ Mr.DETR 4 Route-1 49.9 67.4 54.7 31.1 53.2 65.3
w/ Mr.DETR 4 Route-3 50.0 67.5 54.7 31.3 53.3 65.7
Baseline [70] 5 - 12 46.5 64.2 50.8 28.8 50.0 60.7
w/ Mr.DETR 5 Route-2 49.5 (+3.0) 66.6 54.1 30.3 52.6 64.7
w/ Mr.DETR 5 Route-1 49.8 67.3 54.6 31.0 53.0 65.3
w/ Mr.DETR 5 Route-3 49.9 67.3 54.6 30.8 53.2 65.4
Baseline [70] 0 - 24 41.3 56.3 45.4 25.5 44.6 53.2
w/ Mr.DETR 0 Route-2 42.6 (+1.3) 57.3 46.4 25.8 44.8 55.9
w/ Mr.DETR 0 Route-1 48.2 65.8 52.7 29.1 51.4 62.5
w/ Mr.DETR 0 Route-3 48.2 66.0 52.8 29.2 51.6 62.2
Baseline [70] 1 - 24 45.1 61.5 49.5 28.2 48.3 58.1
w/ Mr.DETR 1 Route-2 47.1 (+2.0) 63.3 51.3 28.4 49.9 61.2
w/ Mr.DETR 1 Route-1 49.6 67.1 54.2 30.5 52.9 63.9
w/ Mr.DETR 1 Route-3 49.7 67.3 54.2 30.6 53.0 64.1
Baseline [70] 2 - 24 47.2 64.3 51.8 29.8 50.3 60.8
w/ Mr.DETR 2 Route-2 49.2 (+2.0) 66.2 53.6 30.2 52.2 63.4
w/ Mr.DETR 2 Route-1 50.1 67.7 54.8 31.1 53.3 64.8
w/ Mr.DETR 2 Route-3 50.2 67.9 54.7 31.4 53.5 64.7
Baseline [70] 3 - 24 48.1 65.7 52.7 30.8 51.3 61.8
w/ Mr.DETR 3 Route-2 50.0 (+1.9) 67.4 54.4 30.9 53.0 64.5
w/ Mr.DETR 3 Route-1 50.2 68.0 54.8 31.7 53.4 64.8
w/ Mr.DETR 3 Route-3 50.3 68.1 54.8 31.5 53.4 65.0
Baseline [70] 4 - 24 48.6 66.3 53.1 30.9 51.8 62.3
w/ Mr.DETR 4 Route-2 50.3 (+1.7) 67.9 54.7 31.3 53.2 65.0
w/ Mr.DETR 4 Route-1 50.5 68.3 55.1 31.7 53.7 65.1
w/ Mr.DETR 4 Route-3 50.4 68.3 54.9 32.0 53.5 65.3
Baseline [70] 5 - 24 48.6 66.4 53.1 30.8 51.9 62.9
w/ Mr.DETR 5 Route-2 50.3 (+1.7) 68.0 54.7 31.5 53.2 65.0
w/ Mr.DETR 5 Route-1 50.4 68.2 55.1 31.5 53.5 65.2
w/ Mr.DETR 5 Route-3 50.4 68.2 54.8 31.7 53.4 65.2
Table 11: Evaluation results of box prediction in all six decoder layers. ‘Route-2’: the primary route for one-to-one prediction. ‘Route-1’: the auxiliary route with an independent FFN. ‘Route-3’: the auxiliary route with an instructive self-attention.
Appendix H Qualitative Results
We present the prediction results of our method in Fig. 10. The model is based on the DINO [61] with IA-BCE loss [1] as the baseline using Swin-L [38] backbone.
Figure 10: Qualitative results of our method. Left: prediction results. Right: ground truth.

### v4

@@ IV-C Instance Segmentation
We integrate our method with the detection transformer architecture for the instance segmentation task. Specifically, following [11, 14, 96], we develop an object mask prediction head based on Deformable-DETR++ [11] utilizing 300 queries. For simplicity, both one-to-one and one-to-many assignments consider only box and classification costs, consistent with the detection transformer. Also, for the IoU score learned in Mr. DETR++, we only consider the box IoU as the localization quality measure. We present the AP metric for both box and mask predictions in Tab. VI. Over a training schedule of 12 and 24 epochs, Mr. DETR enhances the mask AP by approximately 3.6% and 2.5% compared to the baseline model, respectively. Mr. DETR++ can improve 5.3% and 3.9% mask AP over Deformable-DETR++ [11]. These experimental results highlight the effectiveness of our method.
Convergency Curves. Employing Deformable-DETR++ [11] with 300 queries, we perform training on the instance segmentation task both with and without the integration of our proposed approach. Consistent with established methods [11, 9, 14, 13, 8], models are trained using 12 and 24 epoch schedules, respectively. The learning rate is reduced by a factor of 0.1 at the 11th and 20th epochs according to the 12 and 24 epoch schedules, respectively. We illustrate the evaluation results for bounding box predictions in Fig. 7(a) and for instance mask predictions in Fig. 7(b). The evaluation results demonstrate that our approach significantly enhances the training process of the baseline model.
@@
IV-D Panoptic Segmentation
To further evaluate the versatility of Mr. DETR and Mr. DETR++, we extend their application to the panoptic segmentation task, which requires simultaneous prediction of bounding boxes and instance masks for ‘thing’ classes, as well as semantic masks for ‘stuff’ classes. We adopt Mask DINO [97] as our baseline model. To accommodate panoptic segmentation, we modify the one-to-many label assignment strategy in Eqn. (2), as IoU-based matching is unsuitable for semantic mask predictions. Specifically, we implement a one-to-many assignment for ‘stuff’ classes by repeating target labels three times, determined empirically for simplicity. Additionally, the scoring calibration is not used in panoptic segmentation. As reported in Tab. VII, Mr. DETR achieves a 1.2% PQ improvement over the Mask DINO baseline, while Mr. DETR++ enhances performance by 2.2% PQ. Additionally, we provide mask AP and box AP for ‘thing’ classes, highlighting the superior performance of our approach.
TABLE VII: Panoptic segmentation results on the COCO 2017 [90] validation set. All experiments are based on ResNet-50 [91] as the backbone.
Model Epochs Queries PQ PQth PQst APbox APmask
DETR [1] 500 100 43.4 48.2 36.3 - 31.1
Mask2Former [98] 12 100 46.9 52.5 38.4 - 37.2
Panoptic Segformer [99] 12 353 48.0 52.3 41.5 - -
Mask DINO [97] 12 300 49.0 54.8 40.2 43.2 40.4
Mask DINO [97] w/ Mr. DETR 12 300 50.2 (+1.2) 56.6 (+1.8) 40.4 (+0.2) 45.4 (+2.2) 42.0 (+1.6)
Mask DINO [97] w/ Mr. DETR++ 12 300 51.2 (+2.2) 57.5 (+2.7) 41.7 (+1.5) 46.0 (+2.8) 42.5 (+2.1)
IV-E Effectiveness Analysis
@@
IV-F Ablation Study
Effectiveness of Multi-route Training in Mr. DETR. Our method features a primary route, referred to as Route-2 in Fig. 4, dedicated to one-to-one prediction, and two auxiliary routes, Route-1 and Route-3 for one-to-many prediction. Route-1 utilizes shared self-attention and cross-attention with Route-2, but employs an independent FFN to achieve one-to-many prediction. We introduce an instructive self-attention mechanism in Route-3, which shares all parameters with the primary route. To evaluate the contribution of each auxiliary route, we perform an ablation study, as shown in Tab. VIII. Route-1, with its independent FFN, enhances the baseline model by 2.0% mAP. Experimental results indicate that Route-3 provides a 2.8% improvement on Deformable-DETR++ [11], highlighting the effectiveness of our proposed instructive self-attention. By incorporating multi-route training, Deformable-DETR++ [11] achieves 50.7% mAP, surpassing the baseline by 3.1%.
Effectiveness of Mr. DETR++. As shown in Table VIII, Mr. DETR++ replaces the auxiliary training route with two independent FFNs integrated with a route-aware MoE, yielding a 0.3% improvement in AP. Incorporating a scale-aware MoE into the encoder further enhances performance by 0.4% AP. Additionally, applying localization-aware score calibration results in a final AP of 51.8%. We investigate the impact of the factor $\phi$ in App. -E. In App. -F, we further demonstrate our model’s performance across different decoder layers and routes, illustrating the effectiveness of our approach.
TABLE VIII: The ablation study of different routes in our method. ‘Route-1 (FFN)’: the auxiliary training route with independent FFN. ‘Route-2’: the primary route for one-to-one prediction. ‘Route-3’: the auxiliary training route with instructive self-attention. ‘Route-1 (MoE)’: the auxiliary training route with route-aware MoE. ‘MoE-Encoder’: the scale-aware MoE in the encoder. ‘Calibration’: the localization-aware score calibration used during inference.
Variants Route-1 (FFN) Route-2 Route-3 Route-1 (MoE) MoE-Encoder Calibration AP AP50 AP75
@@ IV-F Ablation Study
Mr. DETR 50.7 (+3.1) 68.2 55.4
51.0 (+3.4) 68.7 56.1
51.3 (+3.7) 68.5 55.9
51.4 (+3.8) 68.8 56.3
Mr. DETR++ 51.8 (+4.2) 68.9 56.7
@@
Designs of the Instruction Mechanism. To further reduce the trainable parameters in Route-3 and enhance parameter sharing with the primary route, we introduce the instruction mechanism. In Tab. IX, we evaluate the performance of different designs of the instruction mechanism, as discussed in Sec. III-C. Initially, we evaluate two conventional variants: (a) one with independent self-attention and (b) another without self-attention in Route-3. Experimental results indicate that our approach surpasses these variants by 0.7% and 0.8% AP, respectively. Additionally, we implement a variant (c) using independent object queries in Route-3 while sharing all other parameters with the primary route. Our method exceeds this design by 0.8% AP. We also investigate an implementation (d) that integrates instruction tokens into the object queries by addition, unlike our method, which appends them. Experimental results demonstrate the superiority of our approach, which dynamically and flexibly guides object queries to achieve one-to-many predictions.
TABLE IX: Different designs of instructive self-attention.
Variants AP AP50 AP75
(a) Independent self-attention 50.0 67.7 54.8
(b) Remove self-attention 49.9 67.3 54.7
(c) Independent object queries 49.9 67.5 54.6
(d) Instruction by addition 50.1 67.6 54.7
(e) Instruction by concatenation (ours) 50.7 68.2 55.4
TABLE X: The ablation study on the configurations of the route-aware MoE. These experiments evaluate the selected scales within the scale-aware MoE, with downsampling ratios of x64, x32, x16, and x8.
x64 x32 x16 x8 AP AP50 AP75 GFlops
51.3 68.5 55.9 424.1
51.8 68.9 56.7 486.8
51.7 68.7 56.4 503.4
51.9 69.0 56.7 569.4
51.5 68.6 56.2 832.5
Ablation on the Scale-aware MoE. In Mr. DETR++, we extend the MoE to the transformer encoder to enhance feature representations. However, as the transformer encoder processes a combined input sequence of image tokens from multi-scale features, directly applying MoE incurs significant computing costs. To mitigate this, we propose a scale-aware MoE, which assigns an MoE block to low-scale tokens while sharing a single expert across all tokens. To investigate the impact of selected low-scale tokens, we conduct an ablation study on the scales with an activated MoE block. The results, reported in Tab. X, demonstrate that applying the MoE block to x64 downsampled features yields a 0.5% AP improvement. While using MoE on x64, x32, and x16 downsampled features achieves the highest performance of 51.9% AP, it substantially increases computational costs. Thus, we choose to apply MoE on x64 downsampled features, striking an optimal balance between computing efficiency and performance.
[…]
(a) Layer 3 in the Decoder
[…]
(b) Layer 4 in the Decoder
[…]
(c) Layer 5 in the Decoder
Figure 9: Distribution of Activated Experts in the Route-aware MoE. ‘Route-2’: the primary route for one-to-one prediction. ‘Route-3’: the auxiliary training route with instructive self-attention. ‘Route-1’: the auxiliary training route with independent gating.
Distribution of Activated Experts in the Route-aware MoE. To save computing cost, we use a route-aware MoE in the last three layers of the transformer decoder. As shown in Fig. 9, we visualize the distribution of activated experts in each gating function. In layers 3 and 4, overall, all routes tend to prefer three experts, but the two auxiliary training routes exhibit preference patterns distinct from Route-2. For instance, in layer 3, Route-2’s gating shows a high activation for Expert-0, whereas Route-3 and Route-1 do not. However, Route-2’s expert selection still significantly overlaps with that of the other routes; for example, Expert-2 is consistently favored across all three routes. In the final layer, Route-2 demonstrates a strong preference for the first two experts, while the other two routes distribute their preferences more evenly across the first three experts. This suggests that a route-aware MoE can both avoid task conflicts through expert selection and facilitate knowledge sharing.
Training Cost. We measure the training time of different methods in Tab. XI. The training time denotes the average duration of each epoch. We evaluate training time on 8 NVIDIA 3090 GPUs with a batch size of 16. The experimental results indicate that our proposed method achieves an effective trade-off between performance and training costs.
TABLE XI: Comparison of training time (minutes) of various methods. All methods utilize Deformable-DETR++ with 300 queries as the baseline. The training time represents the average duration (minutes) per training epoch.
Model Duration AP
Deformable-DETR++ [11] 84 47.0
H-DETR [8] 104 (+20) 48.7 (+1.7)
MS-DETR [14] 96 (+12) 48.8 (+1.8)
Mr. DETR (ours) 101 (+17) 49.5 (+2.5)
Mr. DETR++ (ours) 127 (+43) 51.0 (+4.0)
V Mechanism Analysis
Does the non-shared component or instructive self-attention disrupt the necessary information for one-to-one prediction in object queries? We observe that the one-to-many route equipped with a non-shared component can still perform well, as discussed in Sec. I. To investigate whether Route-1 (FFN) and Route-3 (instructive self-attention) achieve one-to-many prediction by discarding essential information for one-to-one prediction in object queries, we conduct probing experiments using Mr. DETR as an example. Specifically, we freeze all parameters of the pre-trained Mr. DETR model and attach a trainable classification layer with random initialization to each route. These trainable layers are then trained under a one-to-one assignment to extract any remaining information relevant to one-to-one prediction from the object queries. The results of the probing experiment, as shown in Tab. XII, indicate that the probe layer for Route-2 achieves comparable performance (50.3 AP) to the oracle one-to-one prediction. In contrast, the probe layers for Route-1 and Route-3 perform significantly worse, both reaching only about 20.0 AP, far below the performance when equipping NMS (49.0 AP). This suggests that while the probe layer for Route-2 can extract sufficient information for one-to-one prediction, the probe layers for Route-1 and Route-3 fail to do so. In summary, the experimental results demonstrate that the non-shared component and instructive self-attention significantly disrupt the information necessary for one-to-one prediction in object queries.
TABLE XII: Probing Analysis. ‘Route-2’: the primary route for one-to-one prediction. ‘Route-1’: the auxiliary training route with independent FFN. ‘Route-3’: the auxiliary training route with the instructive self-attention.
Model w/o NMS w/ NMS
Deformable-DETR++ [11] 47.6 -
Mr. DETR Route-2 50.7 50.8
Mr. DETR Route-2 + Probing 50.3 50.5
Mr. DETR Route-1 12.8 51.0
Mr. DETR Route-1 + Probing 19.3 49.0
Mr. DETR Route-3 12.8 51.1
Mr. DETR Route-3 + Probing 20.8 48.9
Training with o2m only 13.3 50.2
+ Probing for o2o 13.3 50.1
Can any non-shared component in the decoder avoid conflict between one-to-one and one-to-many prediction? As discussed in Sec. I, we find that any non-shared component can mitigate the conflict between one-to-one and one-to-many prediction while benefiting the primary one-to-one prediction. To further investigate the mechanism underlying multi-route training, we conduct an analysis experiment, as shown in Tab. XII. Specifically, we first train an oracle model using only one-to-many supervision and then perform probing analysis on the pre-trained model. The experimental results reveal that object queries supervised solely by one-to-many assignment lack the necessary information for one-to-one prediction. In conjunction with the earlier conclusion that non-shared components disrupt the information required for one-to-one prediction, we infer that routes trained without one-to-one supervision cannot retain the information essential for one-to-one prediction. Moreover, any non-shared component inherently disrupts this information, thereby preventing conflict between one-to-one and one-to-many predictions. In summary, these non-shared components play a key role in avoiding conflicts between the two types of prediction.
VI Conclusion
@@
Regarding the model with auxiliary one-to-many training as a multi-task framework, we examine the roles of each component in the transformer decoder for two training targets. Our empirical findings indicate that any independent component in the decoder can effectively learn both one-to-one and one-to-many training targets, even when other components are shared. Building on this insight, we propose a multi-route training mechanism, featuring a primary route and two auxiliary training routes. The first auxiliary route incorporates our proposed instructive self-attention, which dynamically and flexibly guides object queries for one-to-many prediction. The second auxiliary route is enhanced by our proposed route-aware mixture-of-experts, enabling knowledge sharing while mitigating potential conflicts between routes. Additionally, the transformer encoder is further enhanced with our proposed scale-aware MoE, applied to low-scale features to reduce computational cost. Notably, the auxiliary training routes are discarded during inference. Extensive experiments across multiple tasks, including object detection, instance segmentation, and panoptic segmentation, validate the effectiveness and versatility of our approach.